# DeWave: Discrete EEG Waves Encoding for Brain Dynamics to Text Translation

### v1

@@ Appendix B Implementation Details
We release our implementation code through GitHub to contribute to this area. Currently the basic code are available through an anonymous link [^6] https://anonymous.4open.science/r/DeWave-EEG-Translation-6C2E For word-level EEG features, we use the $56$ tokens each with an $840$ embedding size. The codex encoder for word-level features is a 6 layers transformer encoder with head number 8, hidden embedding 512. For raw EEG waves, we clip or pad the EEG waves up to sample point 5500 with a constant value of zero, which scales up to 11 seconds according to the sampling rate of 500 Hz. The codex encoder for raw EEG wave features is illustrated in Section 3.3, where a 6-layer CNN encoder slides through the whole wave and get the embedding sequence. A transformer layer with head number 8 and a $1\times 1$ convolutional layer are combined to fuse multiple EEG channels into one embedding with size 512.
The codex encoder shares the same structure with word-level features. DeWave uses a codex with size 2048 where each codex latent is an embedding with size 512. The ablation study gives a discussion about the codex size. All models are trained on Nvidia V100 and A100 GPUs. For the self-supervised decoding for raw waves, we use a learning rate of 5e-4, and a VQ coefficient of 0.25 for training 35 epochs. For training the codex (stage 1), DeWave uses a learning rate of 5e-4 for 35 epochs. For finetuning the translation (stage 2), DeWave uses a learning rate of 5e-6 for 30 epochs. We use the SGD as the optimizer for training all the models.
@@ Appendix D Codex Visualization
Since the logic is to learn a discrete codex from brain dynamics, it naturally arouse whether the codex value distribution between raw waves and frequency features has differences. In that case, we conduct additional experiments to visualize the learned codex book with T-SNE methods and report the results in Fig 7. Ideally, the purpose of the discrete codex is to make the language model have a better understanding of brain encoding. In that case, the learned codex regardless of whether it is for frequency features or for raw waves, should be approximately the same as the word2vec embedding. In another word, the distribution of the codex should be similar. Fig. 7 supports our expectation, where the codex learned from frequency and raw waves have very similar distributions. However, the frequency codex has more coverage with corner cases and boundaries. We think this is rational since frequency features are naturally easier to distinguish as it has introduced manually selected features. The performance gap between raw waves and frequency features supports this point as well. Still, the similarity of the distribution illustrates the rationality of the learned codex.
Appendix E Generated Samples
In this section, we visualize the generated decoding text results on brain waves and compare them with the ground truth in Table 5 and Table 6. It suggests that the results are even better on long and simple sentences. For example below, even for the ground truth with long and logic as below, the prediction could still match key information throughout the whole sentence.
@@ Appendix E Generated Samples
  People is feasible to guess the meaning of a human based on the translation from brain waves. In the example above, the model recognizes through waves that the University of xxx at Austin. Although it has a factual mistake that the University of California is not at Austin, it still suggests that the model could approximately capture the semantic meaning through non-invasive brain waves. A similar observation applies that the model recognizes that it is the xx American Studies in xx years however the model predicts history American Studies in 1975 rather than the ground truth is Latin American Studies in 1973. Surprisingly, even the years have correlations at this stage.
@@
Appendix F Subject Wise Evaluation
@@
Section 4.3 introduces subject-wise metric evaluation on word-level EEG features by removing the subject to be tested from the training data, then training the model from scratch for testing. The results are shown in Fig 4 in the main paper, where different subjects share the same reading article. However, in supplementary details, we conduct a more detailed subject-wise evaluation in that we respectively train the model on every single subject on task v2.0 and test every subject to report the cross-subject performance. The single subject denotes the model only trained on a single subject on the task v2.0 dataset. For each subject, however, due to limited data, we add all data from task 1.0 as an assistance base. The results for each subject are reported as below.
@@
Here we seleceted subject YAC (Table 7), YAK (Table 8), YDG (Table 9), YFS (Table 10), YSL (Table 11), and YMD (Table 12) to report the performance. We visualize the metrics by clustering the same metrics value of different subjects in one radar chart. It is observed that the model performance might not be optimized if we train and test on the same subject. For example, if we train the model with task2.0 data from subject YFS and test on all subjects, the YFS subject only reaches BLUE-$\{1-4\}$ 43.32, 26.30, 14.68, and 7.62, which is lower than YDS, YAG, YRP, .. etc. which reach 43.79, 26.46, 14.94, and 8.03. This suggests the cross-subject robusty of the proposed DeWave model. Also, since we use the exact same visualization scale for each radar chart, the area of the chart suggests the performance level. It is observed that if we change the training data from subject to subject, the average performance of every subject is affected by a similar trend. We think this phenomenon is caused by the different signal-to-noise ratios. Some subjects might naturally have less noise interference which is easier for the model to learn meaningful features during the training process.
Appendix G Supplementatary Conclusion
@@
In this supplementary material, we give implementation details, training schema, and most importantly, more generated results and subject-wise evaluation of the proposed DeWave model. The generated results suggest a surprisingly good correlation between the model output on brain waves and the ground truth, even in long sentences with logic. Although there do exist factual and cognitive mistakes in the translation, it is still feasible to guess the meaning of a human based on the translation from brain waves. The subject-wise evaluation suggests the DeWave model is stable across different human subjects. Please refer to the tables attached below.
-eps-converted-to.pdf Subject YDG YAG YRP YLS YFS YMD YRH YFR YTL YAC YSL YAK YMS YSD YHS YDR YRK YIS BLEU-1 43.14 43.14 43.14 43.14 42.88 43.14 43.14 41.98 43.14 42.85 43.20 43.14 43.14 43.14 43.14 43.27 43.14 43.14 BLEU-2 26.58 26.58 26.58 26.58 26.89 26.58 26.58 25.31 26.58 26.07 26.78 26.58 26.58 26.58 26.58 26.71 26.58 26.58 BLEU-3 15.67 15.67 15.67 15.67 16.03 15.67 15.67 14.68 15.67 15.30 15.82 15.67 15.67 15.67 15.67 15.83 15.67 15.67 BLEU-4 9.23 9.23 9.23 9.23 9.54 9.23 9.23 8.20 9.23 8.94 9.38 9.23 9.23 9.23 9.23 9.40 9.23 9.23 ROUGE-R 32.81 32.81 32.81 32.81 32.38 32.81 32.81 31.58 32.81 32.53 33.12 32.81 32.81 32.81 32.81 32.89 32.81 32.81 ROUGE-P 41.03 41.03 41.03 41.03 40.91 41.03 41.03 38.81 41.03 40.58 41.39 41.03 41.03 41.03 41.03 40.86 41.03 41.03 ROUGE-F 36.39 36.39 36.39 36.39 36.09 36.39 36.39 34.77 36.39 36.05 36.72 36.39 36.39 36.39 36.39 36.38 36.39 36.39

### v3

@@ Appendix B Implementation Details
We release our implementation code through GitHub to contribute to this area. Currently the basic code are available through an anonymous link [^6] https://github.com/duanyiqun/DeWave For word-level EEG features, we use the $56$ tokens each with an $840$ embedding size. The codex encoder for word-level features is a 6-layer transformer encoder with head number 8, hidden embedding 512. For raw EEG waves, we clip or pad the EEG waves up to sample point 5500 with a constant value of zero, which scales up to 11 seconds according to the sampling rate of 500 Hz. The codex encoder for raw EEG wave features is illustrated in Section 3.3, where a 6-layer CNN encoder slides through the whole wave and gets the embedding sequence. A transformer layer with head number 8 and a $1\times 1$ convolutional layer are combined to fuse multiple EEG channels into one embedding with size 512.
The codex encoder shares the same structure with word-level features. DeWave uses a codex with size 2048 where each codex latent is an embedding with size 512. The ablation study gives a discussion about the codex size. All models are trained on Nvidia V100 and A100 GPUs. For the self-supervised decoding for raw waves, we use a learning rate of 5e-4 and a VQ coefficient of 0.25 for training 35 epochs. For training the codex (stage 1), DeWave uses a learning rate of 5e-4 for 35 epochs. For finetuning the translation (stage 2), DeWave uses a learning rate of 5e-6 for 30 epochs. We use the SGD as the optimizer for training all the models.
@@ Appendix D Codex Visualization
Since the logic is to learn a discrete codex from brain dynamics, it naturally arises whether the codex value distribution between raw waves and frequency features has differences. In that case, we conduct additional experiments to visualize the learned codex book with T-SNE methods and report the results in Fig 7. Ideally, the purpose of the discrete codex is to make the language model have a better understanding of brain encoding. In that case, the learned codex regardless of whether it is for frequency features or raw waves, should be approximately the same as the word2vec embedding. In other words, the distribution of the codex should be similar. Fig. 7 supports our expectation, where the codex learned from frequency and raw waves have very similar distributions. However, the frequency codex has more coverage with corner cases and boundaries. We think this is rational since frequency features are naturally easier to distinguish as it has introduced manually selected features. The performance gap between raw waves and frequency features supports this point as well. Still, the similarity of the distribution illustrates the rationality of the learned codex.
Appendix E Motivation and Preliminary Tests with LLMs
We provide additional insights into our experiments with larger language models. While our primary experiments utilized BART to ensure consistency in the decoder scale with prior works, it was paramount for us to ascertain that the observed improvements emanated from discrete coding and not just a more sophisticated decoder.
The potential of bridging brain activities with larger language models (LLMs) and advancing towards AGI is a significant research avenue. Recognizing this, we recently undertook an ablation study, where we replaced the BART decoder with OPT and Llama V1. Contrary to our expectations, the performance enhancement was modest. Given the vast implications of this area, we previously refrained from including these findings in the main manuscript for reasons of prudence.
Limited by our computational resources, we employed the PyTorch FSDP mode to fine-tune the OPT-1.3B and Llama-1 7B models with half-precision across three epochs. Taking cues from Mini-GPT4’s method for handling visual tokens, the tokenized EEG waves were prompted into LLMs. The performance metrics for our experiments with BART, OPT 1.3B, and Llama-1 7B decoders are tabulated below:
Table 5: Performance metrics for different decoders on the ZuCo dataset.
Source Decoder BLEU-1 BLEU-3 ROUGE-R ROUGE-P ROUGE-F
Word-level features DeWave 41.35 13.92 28.82 33.71 30.69
Word-level features DeWave + OPT 1.3B 41.97 14.06 28.98 33.82 30.86
Word-level features DeWave + Llama-1 7B 42.84 15.03 29.42 35.43 32.05
[…]
Raw Waves DeWave + OPT 1.3B 21.31 5.84 22.09 29.94 25.42
Raw Waves DeWave + Llama-1 7B 22.05 6.03 22.45 30.01 26.08
From the table, it’s evident that while the LLMs offer some enhancement, the gains are not as pronounced as one might expect. This underscores the complexity of the problem and the challenges of bridging brain activities with LLMs. This simple experiment provides a deeper dive into our experiments with larger language models. We believe these findings offer additional perspectives and pave the way for more nuanced research in this domain.
Appendix F Generated Samples
In this section, we visualize the generated decoding text results on brain waves and compare them with the ground truth in Table 6 and Table 7. It suggests that the results are even better on long and simple sentences. For example, even for the ground truth with long and logic as below, the prediction could still match key information throughout the whole sentence.
@@ Appendix F Generated Samples
  People is feasible to guess the meaning of a human based on the translation from brain waves. In the example above, the model recognizes through waves that the University of xxx at Austin. Although it is a factual mistake that the University of California is not at Austin, it still suggests that the model could approximately capture the semantic meaning through non-invasive brain waves. A similar observation applies that the model recognizes that it is the xx American Studies in xx years however the model predicts history American Studies in 1975 rather than the ground truth is Latin American Studies in 1973. Surprisingly, even the years have correlations at this stage.
@@
Appendix G Subject Wise Evaluation
@@
Section 4.3 introduces subject-wise metric evaluation on word-level EEG features by removing the subject to be tested from the training data, and then training the model from scratch for testing. The results are shown in Fig 4 in the main paper, where different subjects share the same reading article. However, in supplementary details, we conduct a more detailed subject-wise evaluation in that we respectively train the model on every single subject on task v2.0 and test every subject to report the cross-subject performance. The single subject denotes the model only trained on a single subject on the task v2.0 dataset. For each subject, however, due to limited data, we add all data from task 1.0 as an assistance base. The results for each subject are reported below.
@@
Here we seleceted subject YAC (Table 8), YAK (Table 9), YDG (Table 10), YFS (Table 11), YSL (Table 12), and YMD (Table 13) to report the performance. We visualize the metrics by clustering the same metrics value of different subjects in one radar chart. It is observed that the model performance might not be optimized if we train and test on the same subject. For example, if we train the model with task2.0 data from subject YFS and test on all subjects, the YFS subject only reaches BLUE-$\{1-4\}$ 43.32, 26.30, 14.68, and 7.62, which is lower than YDS, YAG, YRP, .. etc. which reach 43.79, 26.46, 14.94, and 8.03. This suggests the cross-subject robustness of the proposed DeWave model. Also, since we use the same visualization scale for each radar chart, the area of the chart suggests the performance level. It is observed that if we change the training data from subject to subject, the average performance of every subject is affected by a similar trend. We think this phenomenon is caused by the different signal-to-noise ratios. Some subjects might naturally have less noise interference which makes it easier for the model to learn meaningful features during the training process.
Appendix H Supplementatary Conclusion
@@
In this supplementary material, we give implementation details, training schema, and most importantly, more generated results and subject-wise evaluation of the proposed DeWave model. The generated results suggest a surprisingly good correlation between the model output on brain waves and the ground truth, even in long sentences with logic. Although there are factual and cognitive mistakes in the translation, it is still feasible to guess the meaning of a human based on the translation from brain waves. The subject-wise evaluation suggests the DeWave model is stable across different human subjects. Please refer to the tables attached below.
Subject
YDG
YAG
YRP
YLS
YFS
YMD
YRH
YFR
YTL
YAC
YSL
YAK
YMS
YSD
YHS
YDR
YRK
YIS
BLEU-1
43.14
43.14
43.14
43.14
42.88
43.14
43.14
41.98
43.14
42.85
43.20
43.14
43.14
43.14
43.14
43.27
43.14
43.14
BLEU-2
26.58
26.58
26.58
26.58
26.89
26.58
26.58
25.31
26.58
26.07
26.78
26.58
26.58
26.58
26.58
26.71
26.58
26.58
BLEU-3
15.67
15.67
15.67
15.67
16.03
15.67
15.67
14.68
15.67
15.30
15.82
15.67
15.67
15.67
15.67
15.83
15.67
15.67
BLEU-4
9.23
9.23
9.23
9.23
9.54
9.23
9.23
8.20
9.23
8.94
9.38
9.23
9.23
9.23
9.23
9.40
9.23
9.23
ROUGE-R
32.81
32.81
32.81
32.81
32.38
32.81
32.81
31.58
32.81
32.53
33.12
32.81
32.81
32.81
32.81
32.89
32.81
32.81
ROUGE-P
41.03
41.03
41.03
41.03
40.91
41.03
41.03
38.81
41.03
40.58
41.39
41.03
41.03
41.03
41.03
40.86
41.03
41.03
ROUGE-F
36.39
36.39
36.39
36.39
36.09
36.39
36.39
34.77
36.39
36.05
36.72
36.39
36.39
36.39
36.39
36.38
36.39
36.39